25.0cm3 of 0.200mol dm-3 sodium hydroxide reacted with 28.7 cm3 of sulphuric acid, calculate the concentration of the sulphuric acid. with calculations


Thank you


1
Expert's answer
2021-07-01T05:30:56-0400

The equation for the reaction involved is;

"2NaOH+H_2SO_4 \to Na_2SO_4+2H_2O"

Number of moles of NaOH = (25.0/1000)x0.100 = 0.00250 moles of NaOH


find the number of moles of H₂SO₄ are required from to neutralize this NaOH


mole ratio "NaOH:H_2SO_4" is 2:1

therefore number of moles of H₂SO₄ = "0.00250\/2=" 0.00125 moles H₂SO₄


concentration of "H_2SO_4" "=" 0.00125/(28.7/1000) = 0.0436 moldm⁻³